OpenStudy (anonymous):

If line segment A'B'= 10, and it went through a dilation transformation with a scale factor of 1/5, what is line segment AB? Select one: a. 50 b. 20 c. -5 d. 2

jimthompson5910 (jim_thompson5910):

x = length of AB (length of AB)*(scale factor) = length of A'B' x*(1/5) = 10 x = ??

OpenStudy (anonymous):

x=20?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: multiply both sides of x*(1/5) = 10 by 5

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):